How far apart are the centers of two objects with masses of 4100 kg and 6500 kg. The graviational attraction between them is 86 N.


1
Expert's answer
2021-12-20T10:26:59-0500

The gravity force (regardless of the object) is given as follows:



"F = G\dfrac{m_1m_2}{r^2}"

where "G = 6.8 \u00d7 10^{-11} m^3 kg^{-1} c^{-2}""m_1 = 4100kg, m_2 = 6500kg", "r" is the distance between the centers (assuming the objects have spherical shape). Expressing "r" and substituting "F = 86N", obtain:


"r = \sqrt{\dfrac{Gm_1m_2}{F}} = \sqrt{\dfrac{6.8 × 10^{-11} \cdot 4100\cdot 6500}{86}} \approx 4.6\times 10^3m"

Answer. 4600 km.